# Rocky planet formation in compact disks around M dwarfs

### v2

@@ Appendix A Gas disk model
We set a disk lifetime of , as very low mass stars could retain their primordial disks for longer times than more massive stars, up to tens of Myr (e.g. Damjanov et al. 2007; Bayo et al. 2012; Downes et al. 2015; Manzo-Martínez et al. 2020). We assumed that the main responsible for gas dispersion is the accretion onto the star. We ignored the effect of the disk photoevaporation produced by the radiation from the star, as photoevaporation rates of gas disk around VLMS are estimated to be as low as  (Herczeg 2007), which is lower than the gas accretion rates onto the star during the disk lifetime.
@@
In all the scenarios that we propose, the initial parameters were set at 1 Myr. We fixed the value  thought the disk lifetime. We assumed that the inner edge of the disk is located at the same distance as the co-rotation radius (radius where the stellar rotational period is equal to the orbital rotation period). Thus, for a star of 0.1 M⊙,
considering a initial stellar rotation period  days (Bolmont et al. 2011; Scholz et al. 2018), the inner edge of the disk is  au. In the case of the stars of 0.3 M⊙ and 0.5 M⊙, assuming  days (Bouvier et al. 2014), the inner edge of the disks are  au and  au, respectively.
@@ B.1 Planet-disk interactions
To estimate the reduction factors for non-circular nor co-planar orbits, as well for the estimation of the acceleration that the embryos or planets suffered due to the gas disk torques, we used the new prescription from Ida et al. (2020). The authors studied the gravitational interactions between the gas and the embryos on the basis of dynamical friction resulting in reduction factors given by:
@@
where , and are versors in the respective directions. The gas velocity is given by , with the Keplerian velocity. The , and represent, respectively, the damping timescales of the semi-major axis , eccentricity and inclination of the embryo’s orbit. Considering that the embryo migration due to its interaction with the gas is non-isothermal and assuming the condition , the damping timescales can be expressed as follow
@@
is the timescale from Papaloizou & Larwood (2000) and Tanaka & Ward (2004) in which all physical parameter are evaluated at the semi-major axis of the orbit.
For our standard model, we show in Figure 16 the normalized total toques  associated to a host star of 0.1 M⊙ with , for a range of planetary embryos masses and semi-major axis, assuming  and considering the gas disk model at 1 Myr. It can be seen that except for a narrow area close to the inner edge of the disk and masses lower than 0.3 M⊕ where the total torques are positive, for the rest of the combinations of mass and semi-major axis, the total torque reminds negative. For a planet in a quasi-circular orbit, this means that a negative torque indicates an inward migration of the planet, while a positive torque indicates an outward migration. This pattern in the sign of the total torque remind the same at different times along the disk lifetime. The difference is that the area where the torques are positive is even narrower and closer to the inner edge of the disk for even lower values of planetary masses, until it almost disappears at 10 Myr.
@@ B.2 Star-planet tidal interactions and general relativistic correction
Tidal interactions between the planets (or embryos) and the star are necessary to be taken into account when studying planet formation around very low mass stars as they have an impact in the orbital evolution of close-in planets leading to the precession of their argument of periastron as well as their semi-major axis and eccentricity decays (e.g. Bolmont et al. 2011; Sánchez et al. 2020). It is also relevant to include the evolution of some physical parameters of the star, such as stellar radius and rotational period of the star as they change significantly during the first tens of million years (Baraffe et al. 2015). The star-planet interactions considered in our simulations follow the equilibrium tide model from Hut (1981) and Eggleton et al. (1998). We included tidal distortions and dissipation terms, considering the tide raised by the host star on each embryo or planet and by each embryo or planet on the star and neglected the tide between embryos or planets as follows,
where and are the potential Love numbers of degree 2 of the star and the embryos, respectively. For the star we assumed the Love number of a M dwarf and for the embryos, the Love number estimated for the Earth (Bolmont et al. 2015). The variable r is the position vector of the embryo with respect to the host star, , is the gravitational constant and , , and are, respectively, the masses and radius of the star and each embryo under the approximation that these objects can instantaneously adjust their equilibrium shapes to the tidal force and considering only up to the second-order harmonic distortions (Darwin 1908). The velocity vector of the embryo v, and are the time-lag model constants for the star and each embryo respectively. The factors and are related with the dissipation factors by
@@
We refer the reader to
Sánchez et al. (2020) and Sánchez et al. (2022) for a detailed description of planet-disk interactions as well as tidal and relativistic corrections, and their associated orbital decay timescales.
@@ Appendix C Test simulations
First, we tested the isolation mass that could be reached regarding different semi-major axis. Thus, in Figure 18 we show the planetary mass evolution of four initial seeds of 0.01 M⊕, two of them initially located at  au and the other two at  au. All of them are in circular orbits and evolve by accreting pebbles with different efficiencies (following either Eq. (17) or (18)) associated to each of their locations in a disk of 1 Myr around a star of 0.1 M⊙. Two of the planetary seeds are also affected by planet-disk interactions (see Section B.1), which allow them to migrate inwards at early stages. This makes them reach the isolation masses (see Eq. (24)) faster but with lower values than the ones the other two seeds that were not affected by the interactions with the gas disk could reach. The seed initially located at  au not affected by planet-disk interactions reached an isolation mass of 1 M⊕ at 100,000 yr, while the one that was affected by such interactions reached a slightly smaller mass at 90,000 yr. On the other hand, the seed initially located at  au not affected by the gas disk interactions reached an isolation mass of 3.5 M⊕ at 700.000 yr, while the one that was affected by the interactions reached a much lower value,  M⊕, at 150,000 yr. It can be seen that depending on the initial location of the planets in the disk, the isolation mass that can be reached is different, leading to more massive planets at further distances from the star. Moreover, if planet-disk interactions are taken into account, it will affect the isolation mass and the period of time when it could be reached.
Afterward, we analyze the impact of the eccentricity in the efficiency of pebble accretion. Thus, in Figure 19 the planetary mass evolution by pebble accretion is shown for a planet with initial 0.01 M⊕ located at  au for different eccentricities: 0, 0.001, 0.01 and 0.1. It can be seen that the evolution for  is equivalent to the one of a planet in a circular orbit, but the evolution is enhanced almost by a factor 2 if the planet has an . On the contrary, if the eccentricity of the planet is  the pebble accretion efficiency decay notably, and the planet takes several Myr to reach the isolation mass. This if because in the case of the highest eccentricity proposed in this example, the planet is accreting always in the 3D regime (, see Eq. (20)) while for the lower values of eccentricity, even though at the beginning the planets are accreting in the 3D regime, they quickly switch to the 2D regime (, see Eq. (19)). This shows the importance of the value of the eccentricity in the efficiency of pebble accretion.

### v3

@@ Appendix A Gas disk model
We set a disk lifetime of , as very low-mass stars could retain their primordial disks for longer times than more massive stars, up to tens of Myr (e.g., Damjanov et al. 2007; Bayo et al. 2012; Downes et al. 2015; Manzo-Martínez et al. 2020). We assumed that the main responsible for gas dispersion is the accretion onto the star. We ignored the effect of the disk photoevaporation produced by the radiation from the star, as photoevaporation rates of gas disk around VLMS are estimated to be as low as (Herczeg 2007), which is lower than the gas accretion rates onto the star during the disk lifetime.
@@
In all the scenarios that we propose, the initial parameters were set at 1 Myr. We fixed the value thought the disk lifetime. We assumed that the inner edge of the disk is located at the same distance as the co-rotation radius (radius where the stellar rotational period is equal to the orbital rotation period). Thus, for a star of 0.1 M⊙, considering a initial stellar rotation period days (Bolmont et al. 2011; Scholz et al. 2018), the inner edge of the disk is au. In the case of the stars of 0.3 M⊙ and 0.5 M⊙, assuming days (Bouvier et al. 2014), the inner edge of the disks are au and au, respectively.
@@ B.1 Planet-disk interactions
To estimate the reduction factors for non-circular nor co-planar orbits, as well as for the estimation of the acceleration that the embryos or planets suffered due to the gas disk torques, we used the new prescription from Ida et al. (2020). The authors studied the gravitational interactions between the gas and the embryos on the basis of dynamical friction, resulting in reduction factors given by:
@@
where , and are versors in the respective directions. The gas velocity is given by , with the Keplerian velocity. The , and represent, respectively, the damping timescales of the semi-major axis , eccentricity and inclination of the embryo’s orbit. Considering that the embryo migration due to its interaction with the gas is non-isothermal and assuming the condition , the damping timescales can be expressed as follows
@@
is the timescale from Papaloizou & Larwood (2000) and Tanaka & Ward (2004) in which all physical parameters are evaluated at the semi-major axis of the orbit.
For our standard model, we show in Figure 16 the normalized total toques associated with a host star of 0.1 M⊙ with , for a range of planetary embryos masses and semi-major axis, assuming and considering the gas disk model at 1 Myr. It can be seen that except for a narrow area close to the inner edge of the disk and masses lower than 0.3 M⊕ where the total torques are positive, for the rest of the combinations of mass and semi-major axis, the total torque reminds negative. For a planet in a quasi-circular orbit, this means that a negative torque indicates an inward migration of the planet, while a positive torque indicates an outward migration. This pattern in the sign of the total torque remind the same at different times along the disk lifetime. The difference is that the area where the torques are positive is even narrower and closer to the inner edge of the disk for even lower values of planetary masses, until it almost disappears at 10 Myr.
@@ B.2 Star-planet tidal interactions and general relativistic correction
Tidal interactions between the planets (or embryos) and the star are necessary to be taken into account when studying planet formation around very low-mass stars as they have an impact in the orbital evolution of close-in planets leading to the precession of their argument of periastron as well as their semi-major axis and eccentricity decays (e.g., Bolmont et al. 2011; Sánchez et al. 2020). It is also relevant to include the evolution of some physical parameters of the star, such as stellar radius and rotational period of the star as they change significantly during the first tens of million years (Baraffe et al. 2015). The star-planet interactions considered in our simulations follow the equilibrium tide model from Hut (1981) and Eggleton et al. (1998). We included tidal distortions and dissipation terms, considering the tide raised by the host star on each embryo or planet and by each embryo or planet on the star and neglected the tide between embryos or planets as follows,
where and are the potential Love numbers of degree 2 of the star and the embryos, respectively. For the star we assumed the Love number of a M dwarf and for the embryos, the Love number estimated for the Earth (Bolmont et al. 2015). The variable r is the position vector of the embryo with respect to the host star, , is the gravitational constant and , , and are, respectively, the masses and radius of the star and each embryo under the approximation that these objects can instantaneously adjust their equilibrium shapes to the tidal force and considering only up to the second-order harmonic distortions (Darwin 1908). The velocity vector of the embryo v, and are the time-lag model constants for the star and each embryo, respectively. The factors and are related with the dissipation factors by
@@
We refer to Sánchez et al. (2020) and Sánchez et al. (2022) for a detailed description of planet-disk interactions as well as tidal and relativistic corrections, and their associated orbital decay timescales.
@@ Appendix C Test simulations
First, we tested the isolation mass that could be reached regarding different semi-major axis. Thus, in Figure 18 we show the planetary mass evolution of four initial seeds of 0.01 M⊕, two of them initially located at au and the other two at au. All of them are in circular orbits and evolve by accreting pebbles with different efficiencies (following either Eq. (17) or (18)) associated with each of their locations in a disk of 1 Myr around a star of 0.1 M⊙. Two of the planetary seeds are also affected by planet-disk interactions (see Section B.1), which allow them to migrate inward at early stages. This makes them reach the isolation masses (see Eq. (24)) faster but with lower values than the ones the other two seeds that were not affected by the interactions with the gas disk could reach. The seed initially located at au not affected by planet-disk interactions reached an isolation mass of 1 M⊕ at 100,000 yr, while the one that was affected by such interactions reached a slightly smaller mass at 90,000 yr. On the other hand, the seed initially located at au not affected by the gas disk interactions reached an isolation mass of 3.5 M⊕ at 700.000 yr, while the one that was affected by the interactions reached a much lower value, M⊕, at 150,000 yr. It can be seen that depending on the initial location of the planets in the disk, the isolation mass that can be reached is different, leading to more massive planets at further distances from the star. Moreover, if planet-disk interactions are taken into account, it will affect the isolation mass and the period of time when it could be reached.
Afterward, we analyzed the impact of the eccentricity in the efficiency of pebble accretion. Thus, in Figure 19 the planetary mass evolution by pebble accretion is shown for a planet with initial 0.01 M⊕ located at au for different eccentricities: 0, 0.001, 0.01 and 0.1. It can be seen that the evolution for is equivalent to the one of a planet in a circular orbit, but the evolution is enhanced almost by a factor 2 if the planet has an . On the contrary, if the eccentricity of the planet is the pebble accretion efficiency decay notably, and the planet takes several Myr to reach the isolation mass. This if because in the case of the highest eccentricity proposed in this example, the planet is accreting always in the 3D regime (, see Eq. (20)) while for the lower values of eccentricity, even though at the beginning the planets are accreting in the 3D regime, they quickly switch to the 2D regime (, see Eq. (19)). This shows the importance of the value of the eccentricity in the efficiency of pebble accretion.